OpenStudy (anonymous):

if y is 21 and x is 1.5 what is y when x is 9

OpenStudy (anonymous):

21/1.5 = x/9 cross multiply 1.5(x) = (9)(21) 1.5x = 189 x = 189/1.5 x = 126 check.. 21/1.5 = 126/9 cross multiply 1.5(126) = (9)(21) 189 = 189 (correct) when x is 9, y is 126 any questions ?